OpenStudy (anonymous):

Two negative integers have a sum of -7 and a product of 6. What are the integers? Separate the answers with a comma

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

what two negative numbers can you multiply to get 6?

OpenStudy (anonymous):

(-3)x2=-6

OpenStudy (anonymous):

-3 time -2 = 6. if you add them together will you get -7?

OpenStudy (anonymous):

soooo??

OpenStudy (anonymous):

remember i have to separate with a cooma

OpenStudy (lulubj):

-6,-1

OpenStudy (lulubj):

too easy my friend

OpenStudy (lulubj):

medal?

OpenStudy (mathstudent55):

You can use a system of equations: xy = 6 x + y = -7 x = -7 - y (-7 - y)y = 6 -7y - y^2 = 6 y^2 + 7y + 6 = 0 (y + 1)(y + 6) = 0 y + 1 = 0 or y + 6 = 0 y = -1 or y = -6 xy = 6 or xy = 6 x(-1) = 6 x(-6) = 6 x = -6 x = -1 ANswer: the numbers are -1 and -6